# On Gottschalk’s surjunctivity conjecture for non-uniform cellular automata

### v1

@@ 1.4 Organization of the paper
In Section 2, we prove that the notions of invertibility and stable invertibility are equivalent for NUCA with finite memory (Lemma 1). As a consequence, we relate the surjectivity, the invertibility, and the stable invertibility of a stably injective NUCA in Corollary 2. As another application, we show that pre-injective stably post-surjective NUCA with finite memory must be stably invertible. We then fix in Section 3 the notations and recall the construction of induced local maps of NUCA in Section 3. The proofs of Theorem A, Theorem B, Theorem C, and Theorem D are given respectively in the subsequent Sections 4, 5, 6, and 7.
@@ 2 Invertibility vs stable invertibility
Since $\sigma_{s}$ is invertible, there exist $N\subset G$ finite and $t\in T^{G}$ where $T=A^{A^{N}}$ such that $\sigma_{s}\circ\sigma_{t}=\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. For every $p\in\Sigma(s)$, we obtain from the relation $\sigma_{s}\circ\sigma_{t}=\operatorname{Id}_{A^{G}}$ and [29, Theorem 11.1] some $q\in\Sigma(t)$ such that $\sigma_{p}\circ\sigma_{q}=\operatorname{Id}_{A^{G}}$. Since $q\in\Sigma(t)$ and $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$, another application of [29, Theorem 11.1] shows that there exists $r\in\Sigma(s)$ such that $\sigma_{q}\circ\sigma_{r}=\operatorname{Id}_{A^{G}}$. It follows that $\sigma_{q}$ is invertible and thus $\sigma_{p}\circ\sigma_{q}=\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$. Hence, $\sigma_{s}$ is stably invertible. ∎
@@
It is trivial that (iii)$\implies$(i). The implication (i)$\implies$(ii) results from the definition of invertibility and [29, Theorem A]. Lemma 1 states that (ii)$\implies$(iii) and the proof is thus complete. ∎
@@
The fact that $\tau$ is invertible follows from [29, Theorem 13.4] (see also [6]). We can thus conclude from Lemma 1 that $\tau$ is stably invertible. ∎
@@ 4 Proof of Theorem A
The theorem is trivial if $G$ is finite since every injective selfmap of a finite set is also surjective. Thus, without loss of generality we can suppose that $G$ is infinite. Up to enlarging $M$ if necessary, we can also assume that $1_{G}\in M$. As $s\in S^{G}$ is asymptotically constant, we can find a constant configuration $c\in S^{G}$ and a finite subset $F\subset G$ such that $s|_{G\setminus F}=c|_{G\setminus F}$. Note that as $G$ is infinite, we have $c\in\Sigma(s)$. Since $\sigma_{s}$ is stably injective, we deduce that $\sigma_{c}\colon A^{G}\to A^{G}$ is an injective CA. We infer from the surjunctivity of the group $G$ that $\sigma_{c}$ is also surjective. By Corollary 2 and Remark 1, it follows that $\sigma_{c}$ is invertible. Therefore, there exist a nonempty finite subset $N\subset G$ and a constant configuration $d\in T^{G}$, where $T=A^{A^{N}}$, such that $1_{G}\in N$ and
@@
The set of NUCA with finite memory over the universe $G$ and the alphabet $A$ forms a monoid with respect to the composition operation and we obtain from [29, Theorem 6.2] a configuration $q\in Q^{G}$, where $Q=A^{A^{MN}}$, such that $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$. Note that $MN$ is a memory set of $\sigma_{q}$. As both $\sigma_{s}$ and $\sigma_{d}$ are injective, $\sigma_{q}=\sigma_{s}\circ\sigma_{d}$ is also injective. Since $\sigma_{c}\circ\sigma_{d}=\operatorname{Id}_{A^{G}}$ and $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$ and $s$ is asymptotic to $c$, we deduce that the configuration $q\in Q^{G}$ is asymptotic to the constant configuration $\pi^{G}$ where $\pi\colon A^{MN}\to A^{\{1_{G}\}}$ is the canonical projection $x\mapsto x(1_{G})$. More precisely, $q|_{G\setminus F}=\pi^{G\setminus F}$ since $s|_{G\setminus F}=c|_{G\setminus F}$ so that for every $g\in G\setminus F$ (see (3) and the proof of [29, Theorem 6.2]):
$q(g)=s(g)\circ f^{+N}_{M,g^{-1}d|_{M}}=c(g)\circ f^{+N}_{M,g^{-1}d|_{M}}=\pi$
@@
Let $E=FMN\subset G$ then $E$ is finite and $F\subset E$ as $1\in M\cap N$. It follows that $q|_{G\setminus E}=\pi^{G\setminus E}$. Consider the map $\Phi\colon A^{E}\to A^{E}$ induced by the restriction of $\sigma_{p}$ to $A^{E}$. More specifically, for every $x\in A^{E}$, we define
$\Phi(x)=\sigma_{q}(y)|_{E}$
@@
We conclude that $\sigma_{q}(z)|_{E}=\sigma_{q}(y)|_{E}$ and thus $\Phi$ is well-defined. Observe that
$\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}.$
@@ 5 Proof of Theorem B
The proof follows the same lines, mutatis mutandis, as in the proof of Theorem A. We can suppose that $G$ is infinite since every surjective selfmap of a finite set is also injective. Up to enlarging $M$, we can also assume that $1_{G}\in M$. By hypothesis, there exist a constant configuration $c\in S^{G}$ and a finite subset $F\subset G$ with $s|_{G\setminus F}=c|_{G\setminus F}$. Note again that $c\in\Sigma(s)$ as $G$ is infinite. We infer from the stable post-surjectivity of $\sigma_{s}$ that the CA $\sigma_{c}\colon A^{G}\to A^{G}$ is post-surjective. Since $G$ is post-injunctive, $\sigma_{c}$ is pre-injective and thus invertible by Corollary 3 and Remark 1. Hence, we can find finite subset $N\subset G$ and a constant configuration $d\in T^{G}$, where $T=A^{A^{N}}$, such that $1_{G}\in N$ and
@@
By [29, Theorem 6.2], there exists a configuration $q\in Q^{G}$, where $Q=A^{A^{MN}}$, such that $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$. As $\sigma_{s}$ and $\sigma_{d}$ are surjective, so is $\sigma_{q}=\sigma_{s}\circ\sigma_{d}$. Let $\pi\colon A^{MN}\to A^{\{1_{G}\}}$ be the canonical projection $x\mapsto x(1_{G})$. As $s|_{G\setminus F}=c|_{G\setminus F}$, we have $q|_{G\setminus F}=\pi^{G\setminus F}$ since for every $g\in G\setminus F$ (see the proof of [29, Theorem 6.2]):
@@
where the last equality follows from (5.1). Let $E=FMN\subset G$ then $F\subset E$ as $1\in M\cap N$. It follows that $q|_{G\setminus E}=\pi^{G\setminus E}$. As in the proof of Theorem A, we can write $\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}$ where $\Phi\colon A^{E}\to A^{E}$ is the map given by $\Phi(x)=\sigma_{q}(y)|_{E}$ for every $x\in A^{E}$ and any $y\in A^{G}$ such that $y|_{E}=x$. Then $\Phi$ is surjective because $\sigma_{q}$ and $\operatorname{Id}_{A^{G\setminus E}}$ are both surjective. It follows that $\Phi$ must be injective as $A^{E}$ is finite. Consequently, $\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}$ is injective. Thus by (5.1), we have $\sigma_{s}=\sigma_{q}\circ(\sigma_{d})^{-1}=\sigma_{q}\circ\sigma_{c}$ is injective and thus pre-injective. By the stable post-surjectivity of $\sigma_{s}$ and Corollary 3, we conclude that $\sigma_{s}$ is stably invertible and the proof is thus complete. ∎
@@ 6 Proof of Theorem C
We first prove the following technical lemma which will enable a reduction of Theorem C to Theorem A.
@@
Let $A$ be an alphabet and let $G$ be a group. Let $M\subset G$ be a finite subset and $S=A^{A^{M}}$. Suppose that $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$ for some $s,t\in S^{G}$ and $s$ has uniformly bounded singularity. Then for each $E\subset G$ finite, there exist asymptotically constant configurations $p,q\in S^{G}$ such that $p|_{E}=s|_{E}$, $q|_{E}=t|_{E}$, and $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$.
@@
Up to enlarging $M$, we can assume that $1_{G}\in M=M^{-1}$. Let $E\subset G$ be a finite subset. Up to enlarging $E$, we can suppose without loss of generality that $M\subset E$ and $E=E^{-1}$. As $s$ has uniformly bounded singularity, there exist a constant configuration $c\in A^{A^{M}}$ and a finite subset $F\subset G$ containing $E^{3}$ such that $s|_{FE^{3}\setminus F}=c|_{FE^{3}\setminus F}$. We define an asymptotically constant configuration $p\in S^{G}$ by setting $p|_{FE}=s|_{FE}$ and $p|_{G\setminus FE}=c|_{G\setminus FE}$. We fix $g_{0}\in FE^{2}\setminus FE$ and define $q\in S^{G}$ by $q(g)=t(g_{0})$ if $g\in G\setminus FE$ and $q(g)=t(g)$ if $g\in FE$. Then $q$ is asymptotic to the constant configuration $d\in S^{G}$ defined by $d(g)=t(g_{0})$ for all $g\in G$. Since $1_{G}\in M$, we have a projection $\pi\colon A^{M^{2}}\to A$ given by $x\mapsto x(1_{G})$. Since $E\subset E^{3}\subset F\subset FE$, we deduce from our construction that $p|_{E}=s|_{E}$ and $q|_{E}=t|_{E}$. To conclude, we only need to check that $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$.
@@
Proof of Theorem C. Since $\sigma_{s}$ is stably injective by hypothesis, we deduce from [29, Theorem A] that there exist a finite subset $N\subset G$ and $t\in T^{G}$, where $T=A^{A^{N}}$, such that $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. Up to enlarging $M$ and $N$, we can assume that $1_{G}\in M=N$ and thus $S=T$. By Corollary 2, it suffices to show that $\sigma_{s}$ is surjective to conlude that $\sigma_{s}$ is stably invertible. We suppose on the contrary that $\sigma_{s}$ is not surjective. Since $\Gamma=\sigma_{s}(A^{G})$ is closed in $A^{G}$ with respect to the prodiscrete topology by [29, Theorem 4.4], there must exist a nonempty finite subset $E\subset G$ such that $\Gamma_{E}\subsetneq A^{E}$. Since $s$ has uniformly bounded singularity, we infer from Lemma 3 that there exist asymptotically constant configurations $p,q\in S^{G}$ such that $p|_{E}=s|_{E}$, $q|_{E}=t|_{E}$, and $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$. Let $\Lambda=\sigma_{p}(A^{G})$ then it follows from $p|_{E}=s|_{E}$ that $\Lambda_{E}=\Gamma_{E}\subsetneq A^{E}$. We deduce that $\sigma_{p}$ is not surjective. In particular, $\sigma_{p}$ is not invertible. On the other hand, the condition $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$ shows that $\sigma_{p}$ is stably injective by [29, Theorem A]. We can thus apply Theorem A to deduce that $\sigma_{p}$ is invertible and thus surjective. Therefore, we obtain a contradiction and the proof is complete. ∎

### v2

@@ 1.4 Organization of the paper
In Section 2, we prove that the notions of invertibility and stable invertibility are equivalent for NUCA with finite memory (Lemma 1). As a consequence, we relate the surjectivity, the invertibility, and the stable invertibility of a stably injective NUCA in Corollary 3. As another application, we show that pre-injective stably post-surjective NUCA with finite memory must be stably invertible. We then fix in Section 3 the notations and recall the construction of induced local maps of NUCA in Section 3. The proofs of Theorem A, Theorem B, Theorem C, Theorem D, and Theorem E are given respectively in the subsequent Sections 4, 5, 6, 7, and 8. The last expository Section 9 contains a more self-contained and constructive proof of Theorem B in the case of finitely generated sofic group universes.
@@ 2 Invertibility vs stable invertibility
Since $\sigma_{s}$ is invertible, there exist $N\subset G$ finite and $t\in T^{G}$ where $T=A^{A^{N}}$ such that $\sigma_{s}\circ\sigma_{t}=\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. For every $p\in\Sigma(s)$, we obtain from the relation $\sigma_{s}\circ\sigma_{t}=\operatorname{Id}_{A^{G}}$ and [38, Theorem 11.1] some $q\in\Sigma(t)$ such that $\sigma_{p}\circ\sigma_{q}=\operatorname{Id}_{A^{G}}$. Since $q\in\Sigma(t)$ and $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$, another application of [38, Theorem 11.1] shows that there exists $r\in\Sigma(s)$ such that $\sigma_{q}\circ\sigma_{r}=\operatorname{Id}_{A^{G}}$. It follows that $\sigma_{q}$ is invertible and thus $\sigma_{p}\circ\sigma_{q}=\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$. Hence, $\sigma_{s}$ is stably invertible. ∎
@@
[…]
@@
The fact that $\tau$ is invertible follows from [38, Theorem 13.4] (see also [8]). We can thus conclude from Lemma 1 that $\tau$ is stably invertible. ∎
@@ 4 Proof of Theorem A
We give below the proof of the stable invertibility of stably injective NUCA which are local perturbations of CA over a surjunctive group universe.
@@
Theorem A is trivial if $G$ is finite since every injective selfmap of a finite set is also surjective. Thus, without loss of generality we can suppose that $G$ is infinite. Up to enlarging $M$ if necessary, we can also assume that $1_{G}\in M$. As $s\in S^{G}$ is asymptotically constant, we can find a constant configuration $c\in S^{G}$ and a finite subset $F\subset G$ such that $s|_{G\setminus F}=c|_{G\setminus F}$. Note that as $G$ is infinite, we have $c\in\Sigma(s)$. Since $\sigma_{s}$ is stably injective, we deduce that $\sigma_{c}\colon A^{G}\to A^{G}$ is an injective CA. We infer from the surjunctivity of the group $G$ that $\sigma_{c}$ is also surjective. By Corollary 3 and Remark 1, it follows that $\sigma_{c}$ is invertible. Therefore, there exist a nonempty finite subset $N\subset G$ and a constant configuration $d\in T^{G}$, where $T=A^{A^{N}}$, such that $1_{G}\in N$ and
@@
The set of NUCA with finite memory over the universe $G$ and the alphabet $A$ forms a monoid with respect to the composition operation and we obtain from [38, Theorem 6.2] a configuration $q\in Q^{G}$, where $Q=A^{A^{MN}}$, such that $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$. Note that $MN$ is a memory set of $\sigma_{q}$. As both $\sigma_{s}$ and $\sigma_{d}$ are injective, $\sigma_{q}=\sigma_{s}\circ\sigma_{d}$ is also injective. Since $\sigma_{c}\circ\sigma_{d}=\operatorname{Id}_{A^{G}}$ and $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$ and $s$ is asymptotic to $c$, we deduce that the configuration $q\in Q^{G}$ is asymptotic to the constant configuration $\pi^{G}$ where $\pi\colon A^{MN}\to A^{\{1_{G}\}}$ is the canonical projection $x\mapsto x(1_{G})$. More precisely, $q|_{G\setminus F}=\pi^{G\setminus F}$ since $s|_{G\setminus F}=c|_{G\setminus F}$ so that for every $g\in G\setminus F$ (see (3) and the proof of [38, Theorem 6.2]):
$q(g)=s(g)\circ f^{+N}_{M,g^{-1}d|_{M}}=c(g)\circ f^{+N}_{M,g^{-1}d|_{M}}=\pi$ (4.2)
@@
Let $E=FMN\subset G$ then $E$ is finite and $F\subset E$ as $1_{G}\in M\cap N$. It follows that $q|_{G\setminus E}=\pi^{G\setminus E}$. Consider the map $\Phi\colon A^{E}\to A^{E}$ induced by the restriction of $\sigma_{p}$ to $A^{E}$. More specifically, for every $x\in A^{E}$, we define
$\Phi(x)=\sigma_{q}(y)|_{E}$ (4.3)
@@
We conclude that $\sigma_{q}(z)|_{E}=\sigma_{q}(y)|_{E}$ and thus $\Phi$ is well-defined.
As $MN$ is a memory set of $\sigma_{q}$, we infer from (4.2) and the relation $E=FMN$ that $\sigma_{q}(x)|_{G\setminus E}=x|_{G\setminus E}$ for every $x\in A^{G}$. Combining with (4.3), we obtain
$\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}\colon A^{E}\times A^{G\setminus E}\to A^{E}\times A^{G\setminus E}$
because for every $x=(y,z)\in A^{E}\times A^{G\setminus E}=A^{G}$ where $y=x|_{E}$ and $z=x|_{G\setminus E}$, we have $\sigma_{q}(x)=(\sigma_{q}(x)|_{E},\sigma_{q}(x)|_{G\setminus E})=(\Phi(y),z)\in A^{G}$. We note here that such a map $\sigma_{q}$ is related to the notion of a gate on a subshift [45][Lemma 15].
@@ 5 Proof of Theorem B
By following the same lines, mutatis mutandis, as in the proof of Theorem A, we obtain below the proof of Theorem B on the stable invertibility of stably post-surjective NUCA which are local perturbations of CA over post-injunctive group universes.
@@
We can suppose without loss of generality that $G$ is infinite since every surjective selfmap of a finite set is also injective. Up to enlarging $M$, we can also assume that $1_{G}\in M$. By hypothesis, there exist a constant configuration $c\in S^{G}$ and a finite subset $F\subset G$ with $s|_{G\setminus F}=c|_{G\setminus F}$. Note again that $c\in\Sigma(s)$ as $G$ is infinite. We infer from the stable post-surjectivity of $\sigma_{s}$ that the CA $\sigma_{c}\colon A^{G}\to A^{G}$ is post-surjective. Since $G$ is post-injunctive, $\sigma_{c}$ is pre-injective and thus invertible by Corollary 4 and Remark 1. Hence, we can find finite subset $N\subset G$ and a constant configuration $d\in T^{G}$, where $T=A^{A^{N}}$, such that $1_{G}\in N$ and
@@
By [38, Theorem 6.2], there exists a configuration $q\in Q^{G}$, where $Q=A^{A^{MN}}$, such that $\sigma_{s}\circ\sigma_{d}=\sigma_{q}$. As $\sigma_{s}$ and $\sigma_{d}$ are surjective, so is $\sigma_{q}=\sigma_{s}\circ\sigma_{d}$. Let $\pi\colon A^{MN}\to A^{\{1_{G}\}}$ be the canonical projection $x\mapsto x(1_{G})$. As $s|_{G\setminus F}=c|_{G\setminus F}$, we have $q|_{G\setminus F}=\pi^{G\setminus F}$ since for every $g\in G\setminus F$ (see the proof of [38, Theorem 6.2]):
@@
where the last equality follows from (5.1). Let $E=FMN\subset G$ then $F\subset E$ as $1_{G}\in M\cap N$. It follows that $q|_{G\setminus E}=\pi^{G\setminus E}$. As in the proof of Theorem A, we can write $\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}$ where $\Phi\colon A^{E}\to A^{E}$ is the map given by $\Phi(x)=\sigma_{q}(y)|_{E}$ for every $x\in A^{E}$ and any $y\in A^{G}$ such that $y|_{E}=x$. Then $\Phi$ is surjective because $\sigma_{q}$ and $\operatorname{Id}_{A^{G\setminus E}}$ are both surjective. It follows that $\Phi$ must be injective as $A^{E}$ is finite. Consequently, $\sigma_{q}=\Phi\times\operatorname{Id}_{A^{G\setminus E}}$ is injective. Thus by (5.1), we have $\sigma_{s}=\sigma_{q}\circ(\sigma_{d})^{-1}=\sigma_{q}\circ\sigma_{c}$ is injective and thus pre-injective. By the stable post-surjectivity of $\sigma_{s}$ and Corollary 4, we conclude that $\sigma_{s}$ is stably invertible and the proof is thus complete. ∎
@@ 6 Proof of Theorem C
The goal of this section is to establish Theorem C which states that over surjunctive group universes, every stably injective NUCA with finite memory and uniformly bounded singularity must be stably invertible. We first prove the following useful technical lemma which will enable a reduction of Theorem C to Theorem A.
@@
Let $A$ be an alphabet and let $G$ be a finitely generated group. Let $M\subset G$ be a finite subset and $S=A^{A^{M}}$. Suppose that $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$ for some $s,t\in S^{G}$ and $s$ has uniformly bounded singularity. Then for each $E\subset G$ finite, there exist asymptotically constant configurations $p,q\in S^{G}$ such that $p|_{E}=s|_{E}$, $q|_{E}=t|_{E}$, and $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$.
@@
Up to enlarging $M$, we can assume that $1_{G}\in M=M^{-1}$. Let $E\subset G$ be a finite subset. Up to enlarging $E$, we can suppose without loss of generality that $M\subset E=E^{-1}$ and $E$ is a finite generating set of $G$. As $s$ has uniformly bounded singularity, there exist a constant configuration $c\in S^{G}$ and a finite subset $F\subset G$ containing $E^{3}$ such that $s|_{FE^{3}\setminus F}=c|_{FE^{3}\setminus F}$. We define an asymptotically constant configuration $p\in S^{G}$ by setting $p|_{FE}=s|_{FE}$ and $p|_{G\setminus FE}=c|_{G\setminus FE}$. Note that $FE^{2}\setminus FE\neq\varnothing$. Indeed, if $FE^{2}\subset FE$ then by induction $FE^{n}\subset FE$ for every integer $n\geq 1$ and thus $G=\bigcup_{n\geq 1}E^{n}\subset\bigcup_{n\geq 1}FE^{n}\subset FE$ which is a contradiction since $G$ is infinite. Therefore, we can fix $g_{0}\in FE^{2}\setminus FE$ and define $q\in S^{G}$ by $q(g)=t(g_{0})$ if $g\in G\setminus FE$ and $q(g)=t(g)$ if $g\in FE$. Then $q$ is asymptotic to the constant configuration $d\in S^{G}$ defined by $d(g)=t(g_{0})$ for all $g\in G$. Since $1_{G}\in M$, we have a projection $\pi\colon A^{M^{2}}\to A$ given by $x\mapsto x(1_{G})$. Since $E\subset E^{3}\subset F\subset FE$, we deduce from our construction that $p|_{E}=s|_{E}$ and $q|_{E}=t|_{E}$. To conclude, we only need to check that $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$.
@@
Proof of Theorem C. Since $\sigma_{s}$ is stably injective by hypothesis, we deduce from [38, Theorem A] that there exist a finite subset $N\subset G$ and $t\in T^{G}$, where $T=A^{A^{N}}$, such that $\sigma_{t}\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. Up to enlarging $M$ and $N$, we can assume that $1_{G}\in M=N$ and thus $S=T$. By Corollary 3, it suffices to show that $\sigma_{s}$ is surjective to conlude that $\sigma_{s}$ is stably invertible. We suppose on the contrary that $\sigma_{s}$ is not surjective. Since $\Gamma=\sigma_{s}(A^{G})$ is closed in $A^{G}$ with respect to the prodiscrete topology by [38, Theorem 4.4], there must exist a nonempty finite subset $E\subset G$ such that $\Gamma_{E}\subsetneq A^{E}$. Since $s$ has uniformly bounded singularity, we infer from Lemma 3 that there exist asymptotically constant configurations $p,q\in S^{G}$ such that $p|_{E}=s|_{E}$, $q|_{E}=t|_{E}$, and $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$. Let $\Lambda=\sigma_{p}(A^{G})$. Then it follows from $p|_{E}=s|_{E}$ that $\Lambda_{E}=\Gamma_{E}\subsetneq A^{E}$. We deduce that $\sigma_{p}$ is not surjective. In particular, $\sigma_{p}$ is not invertible. On the other hand, the condition $\sigma_{q}\circ\sigma_{p}=\operatorname{Id}_{A^{G}}$ shows that $\sigma_{p}$ is stably injective by [38, Theorem A]. We can thus apply Theorem A to deduce that $\sigma_{p}$ is invertible and thus surjective. Therefore, we obtain a contradiction and the proof is complete. ∎